# An approximate Kappa generator for particle simulations

### 21 Mar 2026 v2

@@ I Introduction
In kinetic plasma simulations, it is often necessary to generate particle velocity distributions by using random numbers (random variates) at the initialization stage. It is straightforward to produce the Maxwell distribution, because one can employ a random number generator (RNG) for the normal (Gaussian) distribution, which is built-in in numerous programming languages and libraries. One can also implement the Box–Muller method for generating two independent normal variates from two uniform variates. Typically, three independent normal variates per particle are necessary to initialize a Maxwellian plasma. Related to this, one of the authors has developed an inverse sampling method for the 3D Maxwell distribution (Umeda, 2024). The author approximated the cumulative distribution function of the Maxwellian by using an invertible function, and then constructed an inverse transform procedure.
@@ VI Discussion and Summary
In the first and second tests, we find that the proposed method provides a remarkably good approximation. The results are visually indistinguishable from the exact solutions (Fig. 5), as the difference around the most probable speed $v\approx\theta$ is kept small (Fig. 3(b)). In the low-$\kappa$ regime ($\kappa<4$), the proposed method gives a near-perfect approximation, as highlighted in the low $\mathcal{D}$ values (Fig. 4(a)). The Monte Carlo results with $\kappa=3.0$ are indistinguishable from the Kappa distribution in Fig. 6. Extrapolating the results toward the asymptotic value, we will see a difference when $N_{p}\gtrsim 10^{9.5}$. In the high-$\kappa$ regime ($\kappa\gtrsim 5$), the difference between the two distributions emerges in the high-energy part (Fig. 3(b)). Since phase space density in the high-energy part is already very low for high-$\kappa$ cases, the overall properties of the distribution are unlikely to be changed. Figures 4(a) and 4(b) tell us that the discrepancy is most pronounced in the transition case of $\kappa\approx 4.1$. In Fig. 6, the relative entropies deviate when $N_{p}\gtrsim 10^{7.5}$ ($\kappa\gtrsim 5$) and $N_{p}\gtrsim 10^{7}$ ($\kappa\approx 4.1$). Below these particle numbers, the difference between the approximate and exact Kappa distributions is completely hidden by the particle noise. Since today’s PIC simulation typically uses $10^{2-3}$ particles in a grid cell, the proposed method would be accurate enough in many cases. The difference between the two distributions certainly emerges in $D$ beyond these numbers; however, even in the large-number limit, the PSD profile still looks similar to that of the Kappa distribution (Fig. 3(a)), at least on the logarithmic scale. Probably the biggest concern is the total energy density (Fig. 4(b)). The normalized error is within $10^{-3}$ in both low- and high-$\kappa$ regimes, but it increases to $\mathcal{O}(10^{-2.5})$ in the transition case. If the user wants to explore the energy balance at the accuracy of $\mathcal{O}(10^{-2.5})$, the proposed method is not ideal.
@@
The rate ranges from $0.8$ to $0.73$. Considering these, the Pareto method requires $2+2/p_{2}$ variates on CPUs.
@@
The situation is different on GPUs. Inside GPUs, typically 32 threads are controlled by the same instruction. In such a case, one can proceed to the next computational step, only after all 32 threads complete the loop. According to Ridley & Forget (2021), the typical count $\mu$ of the rejection loop is approximated by the mode of the exponentiated geometric distribution (Chacrabaty & Gupta, 2015),
$\mu_{1}(\rho,t)$ $\approx 3-(1-\rho)^{t}-(1-\rho^{2})^{t}$~~ $({\rm for~small~\rho})$ (30)
$\mu_{2}(\rho,t)$ $\approx\dfrac{\log\left(\frac{2(\rho-1)t}{\log(\rho)}\right)}{\log(1/\rho)}+1$ $({\rm for~large~\rho})$ (31)
where $\rho=1-p$ is the rejection probability at each step and $t=32$ is the number of SIMT threads. Using this approximation, from Eqs. (28) and (30), the expected number of variates for the standard method is $3+2\mu_{1}(1-p_{1},32)$ on GPUs. Similarly, from Eqs. (29) and (31), one for the Pareto method is estimated to be $2+2\mu_{2}(1-p_{2},32)$ on GPUs. All these numbers are presented in Figure 8 as a function of $\kappa$.
Although crudely, Figure 8 captures the key trends in Figure 7. The Pareto method becomes substantially slower on GPUs, likely because the rejection loop takes much more iterations on GPUs than on CPUs. Their weak variations on $\kappa$ are also similar. These results further support our argument that these are influenced by the side effects on GPUs. The standard method also runs slower on GPUs, as $\kappa$ approaches $3/2$. We recognize a sharper increase of the elapsed time near $\kappa=1.5$, both in Figure 7(b) and in Figure 8. Probably Figure 8 underestimates the side effect, because the MT method contains some calculations inside the rejection loop. We do not see a slowdown in the proposed method. It behaves similarly regardless of $\kappa$. Since the proposed method does not contain the rejection loop, it does not suffer from the slowdown in SIMT processors. In the future, we may need to use accelerators that employ more SIMT threads than 32. In such a case, the conventional methods perform even worse, while the proposed method should maintain its performance.
@@ Appendix A Taylor expansions
We can also verify this expansion, with help from the hypergeometric function $_{2}F_{1}(a,b;c;x)$ and formulas in Olver et al. (2025, 8.1.7, 15.2.1),
@@
We can similarly check the results using the formulas (Olver et al., 2025, 8.1.7, 15.2.1),
@@
This is in agreement with agreement with Eq. (14).
Next we evaluate the approximate CDF of Eq. (7). We rewrite it by using the $q$-exponential function,
@@
We expand the fraction part,
$\frac{ax+bx^{2}}{1+cx}=ax+(b-ac)x^{2}+\mathcal{O}(x^{3})$ (43)
and then we expand the $q$-exponential function
@@
For $x\rightarrow\infty$, we expand the fraction part with respect to $(1/x)$:
$\frac{ax+bx^{2}}{1+cx}=x\left(\frac{b}{c}+\frac{ac-b}{c^{2}}x^{-1}+\mathcal{O}(x^{-2})\right)$ (46)
Using this, the $q$-exponential function yields

### 09 May 2026 v3

@@ I Introduction
In kinetic plasma simulations, it is often necessary to generate particle velocity distributions by using random numbers (random variates) at the initialization stage. It is straightforward to produce the Maxwell distribution, because one can employ a random number generator (RNG) for the normal (Gaussian) distribution, which is built into numerous programming languages and libraries. One can also implement the Box–Muller method for generating two independent normal variates from two uniform variates (Box and Muller, 1958). Typically, three independent normal variates per particle are necessary to initialize a Maxwellian plasma. Related to this, one of the authors has developed an inverse sampling method for the 3D Maxwell distribution (Umeda, 2024). The author approximated the cumulative distribution function of the Maxwellian by using an invertible function, and then constructed an inverse transform procedure.
@@ VI Discussion and Summary
In the first and second tests, we find that the proposed method provides a remarkably good approximation. The results are visually indistinguishable from the exact solutions (Fig. 5), as the difference around the most probable speed $v\approx\theta$ is kept small (Fig. 3(b)). In the low-$\kappa$ regime ($\kappa<4$), the proposed method gives a near-perfect approximation, as highlighted in the low $\mathcal{D}$ values (Fig. 4(a)). The Monte Carlo results with $\kappa=3.0$ are indistinguishable from the Kappa distribution in Fig. 6. Extrapolating the results toward the asymptotic value, we will see a difference when $N_{p}\gtrsim 10^{9.5}$. In the high-$\kappa$ regime ($\kappa\gtrsim 5$), the difference between the two distributions emerges in the high-energy part (Fig. 3(b)). Since phase space density in the high-energy part is already very low for high-$\kappa$ cases, the overall properties of the distribution are unlikely to be changed. Figures 4(a) and 4(b) tell us that the discrepancy is most pronounced in the transition case of $\kappa\approx 4.1$. In Fig. 6, the relative entropies deviate when $N_{p}\gtrsim 10^{7.5}$ ($\kappa\gtrsim 5$) and $N_{p}\gtrsim 10^{7}$ ($\kappa\approx 4.1$). Below these particle numbers, the difference between the approximate and exact Kappa distributions is completely hidden by the particle noise. Since today’s PIC simulation typically uses $10^{2}$$10^{3}$ particles in a grid cell, the proposed method would be accurate enough in many cases. The difference between the two distributions certainly emerges in $D$ beyond these numbers; however, even in the large-number limit, the PSD profile still looks similar to that of the Kappa distribution (Fig. 3(a)), at least on the logarithmic scale. Probably the biggest concern is the total energy density (Fig. 4(b)). The normalized error is within $10^{-3}$ in both low- and high-$\kappa$ regimes, but it increases to $\mathcal{O}(10^{-2.5})$ in the transition case. If the user wants to explore the energy balance at the accuracy of $\mathcal{O}(10^{-2.5})$, the proposed method is not ideal.
@@
The rate ranges from $0.81$ to $0.73$. Considering these, the Pareto method requires $2+2/p_{2}$ variates on CPUs.
@@
The situation is different on GPUs. Inside GPUs, typically 32 threads are controlled by the same instruction. In such a case, one can proceed to the next computational step, only after all 32 threads complete the loop. According to Ridley & Forget (2021), the typical count $\mu$ of the rejection loop is approximated by the mode of the exponentiated geometric distribution (Chacrabaty & Gupta, 2015). Depending on the rejection rate at each step, $\rho=1-p$, and the number of SIMT threads $t$, $\mu$ is approximated by
$\mu_{1}(\rho,t)$ $\approx 3-(1-\rho)^{t}-(1-\rho^{2})^{t}$~~ $(\rho\lesssim 2^{1/t}-1)$ (30)
$\mu_{2}(\rho,t)$ $\approx\dfrac{\log\left(\frac{2(\rho-1)t}{\log(\rho)}\right)}{\log(1/\rho)}+1$ $(\rho\gg 2^{1/t}-1)$ (31)
For $t=32$, the threshold value $2^{1/t}-1$ is $\approx 0.022$. In the case of the standard method, Eq. (28) gives the rejection rate of $\rho=1-p_{1}(\kappa-1/2)\lesssim 0.05$. This sometimes exceeds the threshold, however, Eq. (30) actually gives a better approximation than Eq. (31) for $\rho\lesssim 0.1$, as shown in Figure 6 in Ridley & Forget (2021). Thus, we estimate the average number of variates to be $3+2\mu_{1}(1-p_{1},32)$ on GPUs. On the other hand, the Pareto method (Eq. (29)) gives the rejection rate of $0.19<(1-p_{2})<0.27$, which is greater than the threshold by an order-of-magnitude. Eq. (31) gives a better approximation in this case. Although it is difficult to see, we can confirm this in Figure 5 in Ridley & Forget (2021). Then the number of variates is estimated to be $2+2\mu_{2}(1-p_{2},32)$ on GPUs. All these numbers are presented in Figure 8 as a function of $\kappa$.
Although crudely, Figure 8 captures the key trends in Figure 7. The Pareto method becomes substantially slower on GPUs, likely because the rejection loop takes much more iterations on GPUs than on CPUs. Their weak variations on $\kappa$ are also similar. These results further support our argument that these are influenced by the side effects on GPUs. The standard method also runs slower on GPUs, as $\kappa$ approaches $3/2$. We recognize a sharper increase of the elapsed time near $\kappa=1.5$, both in Figure 7(b) and in Figure 8. We do not see a slowdown in the proposed method. Its behavior is independent of $\kappa$. Since the proposed method does not contain the rejection loop, it does not suffer from the slowdown in SIMT processors. In the future, we may need to use accelerators that employ more SIMT threads than 32. In such a case, the conventional methods perform even worse, while the proposed method should maintain its performance.
@@ Appendix A Taylor expansions
We can also verify this expansion, with help from the hypergeometric function $_{2}F_{1}(a,b;c;x)$ and formulas in Olver et al. (2026, 8.1.7, 15.2.1),
@@
We can similarly check the results using the formulas (Olver et al., 2026, 8.1.7, 15.2.1),
@@
This is in agreement with Eq. (14).
Next we evaluate the approximate CDF of Eq. (7). We rewrite it as follows,
@@
We expand the $q$-exponential part
@@
For $x\rightarrow\infty$, the $q$-exponential part yields